# An improved perturbative QCD study of the decays $B_{c}^{+}\to\eta_{c}L^{+}$

### v2

@@ III.2 Factorizable-emission suppressed $B_{c}^{+}\to\eta_{c}L^{+}$
In order to understand this peculiar feature in the $B_{c}^{+}\to\eta_{c}S^{+}$ decays, the related decay amplitudes calculated in the iPQCD formalism are presented explicitly in Table 4. Evidently different to the $B_{c}^{+}\to J/\psi S^{+}$ decay amplitudes as shown in Table II of Liu:2023kxr , the originally naive anticipation of constructive interferences between the nonfactorizable-emission diagrams Fig. 1(c) and 1(d) with anti-symmetric leading-twist distribution amplitude of light scalars does not appear. In contrast, the sharply destructive interferences result in much smaller decay amplitudes as exhibited in the third column of Table 4. Furthermore, the destructions in $S1$ are heavier than those in $S2$ for the $B_{c}^{+}\to\eta_{c}(a_{0}(1450),K_{0}^{*}(1430))^{+}$ decays. Because of the allowed $SU(3)$ symmetry breaking effects, the comparably large factorizable-emission contributions, though which are really suppressed to the $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays, lead to further significant destructions between the decay amplitudes of factorizable-emission and nonfactorizable-emission topologies in these $B_{c}^{+}\to\eta_{c}S^{+}$ decays. However, it should be stressed that, due to the light quark masses objectively satisfying the relation, that is, $m_{s}\gg m_{d}\sim m_{u}$, the evidently larger factorizable-emission contributions induced by the vector decay constants $f_{\kappa^{+}},f_{K_{0}^{*}(1430)^{+}}\propto(m_{s}-m_{u})$ Cheng:2005nb are therefore produced in $B_{c}^{+}\to\eta_{c}(\kappa,K_{0}^{*}(1430))^{+}$ channels, relative to the factorizable-emission amplitudes in $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$ modes proportional to the value of $(m_{d}-m_{u})$, namely, tiny broken isospin symmetry.
Nevertheless, the large iPQCD values of ${\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+})$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{0}^{*}(1430)^{+})$ in $S2$ might be tested at the future LHC experiments via resonances $\kappa$ and $K_{0}^{*}(1430)$ with ${\cal B}(\kappa^{+}\to K^{0}\pi^{+})\sim 2/3$ and ${\cal B}(K_{0}^{*}(1430)^{+}\to K^{0}\pi^{+})\sim 0.62$ Maiani:2025hsz under the narrow-width approximation,
${\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+}(\to K^{0}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}\kappa^{+})\cdot{\cal B}(\kappa^{+}\to K^{0}\pi^{+})=(4.80^{+2.46}_{-1.75})\times 10^{-6}\;,$ (91)
and
${\rm BR}(B_{c}^{+}\to\eta_{c}K_{0}^{*}(1430)^{+}(\to K^{0}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}K_{0}^{*}(1430)^{+})\cdot{\cal B}(K_{0}^{*}(1430)^{+}\to K^{0}\pi^{+})$ (92)
$=$ $(8.00^{+5.64}_{-3.60})\times 10^{-6}\;.$
The above iPQCD results predicted for the $B_{c}^{+}\to\eta_{c}S^{+}$ decays await the future experimental tests, which could help us to further explore the complicated QCD dynamics potentially.
@@
As aforementioned, tensor states cannot be produced via the vector current. Therefore, in the $B_{c}^{+}\to\eta_{c}T^{+}$ modes, the factorizable-emission like contributions are forbidden at leading order naturally and their studies must go beyond naive factorization. Then, the $B_{c}^{+}\to\eta_{c}T^{+}$ BRs calculated in the iPQCD formalism are as follows,
@@
associated with the ratios,
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $0.11^{+0.05}_{-0.04}\;,\qquad R^{\eta_{c}}_{a_{2}/\pi}\equiv\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=0.07^{+0.00}_{-0.01}\;.$ (95)
where the dominant errors come from the $B_{c}$-meson shape parameter $\beta_{B_{c}}$.
Based upon the strong decay rate, i.e., ${\cal B}(a_{2}(1320)^{+}\to\pi^{+}\pi^{-}\pi^{+})=0.351\pm 0.135$ ParticleDataGroup:2024cfk , a large BR of $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}(\to\pi^{+}\pi^{-}\pi^{+})$ via resonance $a_{2}(1320)$ could be derived directly under the narrow-width approximation in iPQCD formalism,
${\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}(\to\pi^{+}\pi^{-}\pi^{+}))$ $\equiv$ ${\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})\cdot{\cal B}(a_{2}(1320)^{+}\to\pi^{+}\pi^{-}\pi^{+})$ (96)
$=$ $(4.67^{+2.37}_{-2.20})\times 10^{-5}\;,$
which can be confronted with the near-future examinations in the LHC experiments. The future tests about these values just from nonfactorizable-emission contributions would help us to examine the reliability of this iPQCD formalism.
Similar to $R_{K^{*}/\rho}^{\eta_{c}}$ in the $B_{c}^{+}\to\eta_{c}V^{+}$ sector, another ratio $R^{\eta_{c}}_{K_{2}/a_{2}}$ in the $B_{c}^{+}\to\eta_{c}T^{+}$ decays is also defined by utilizing the $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}$ and $B_{c}^{+}\to\eta_{c}{K_{2}^{*}}(1430)^{+}$ BRs in the iPQCD framework, and its value is then read as,
@@
This result induced by only nonfactorizable-emission contributions is very close to the value presented in Eq. (50), as well as that naively anticipated in factorization ansatz for factorizable-emission predominated $B_{c}^{+}\to\eta_{c}(\rho,K^{*})^{+}$ modes within errors.
@@
Analogous to $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$ but different to $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$, the considerably constructive interferences due to the anti-symmetric leading-twist distribution amplitude between the two nonfactorizable-emission diagrams Fig. 1(c) and 1(d) work in the $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}$ mode indeed. The future measurements on the iPQCD predictions of $B_{c}^{+}\to\eta_{c}L^{+}$ in this work could help test, even differentiate the reliability of the adopted approaches in the literature, which might help us to further understand the rich QCD dynamics in $B_{c}$-meson decays through pulling together disparate ideas.
IV Conclusions and Summary
@@
In summary, we have studied the $B_{c}$-meson decays into $\eta_{c}$ plus a light charged meson in the self-consistent iPQCD formalism at leading order of strong coupling $\alpha_{s}$. The numerical results and phenomenological insights on CP-averaged BRs in association with relative ratios are presented explicitly. The BRs of factorizable-emission dominated decays $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ and $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ are generally larger than those $B_{c}\to J/\psi$ decays correspondingly in the iPQCD formalism, which reveal the remarkably different QCD dynamics between these two kinds of $B_{c}$-meson decays. For (near) future tests at experiments, the BRs of multibody modes of $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ via $\eta_{c}$ resonance through $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$ are also predicted under the narrow-width approximation. The experimental search performed with the successfully upgraded LHCb detector for the BRs around ${\cal O}(10^{-6})$ and above, for example, ${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p})\pi^{+})=(2.70^{+0.74}_{-0.59})\times 10^{-6}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p})\rho^{+})=(7.14^{+2.03}_{-1.60})\times 10^{-6}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}[\to 2(\pi^{+}\pi^{-})]\pi^{+})=(1.95^{+0.59}_{-0.50})\times 10^{-5}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}[\to 2(\pi^{+}\pi^{-})]\rho^{+})=(5.16^{+1.62}_{-1.34})\times 10^{-5}$, $\cdots$, are expected to help probe the relevant decay channels and explore the nature of charmonium $\eta_{c}$. For the factorizable-emission suppressed decays, the almost equal BRs of $B_{c}^{+}\to(J/\psi,\eta_{c})b_{1}(1235)^{+}$ while the surprisingly smaller ones of $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$ than those of $B_{c}^{+}\to J/\psi a_{0}(980,1450)^{+}$ need further explorations in both of theory and experiment. Moreover, the predicted ratios $R_{K/\pi}^{\eta_{c}}$ and $R_{K^{*}/\rho}^{\eta_{c}}$ matching the anticipations based on naive ansatz means that the $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays are predominated by factorizable-emission contributions. The experimental tests on the iPQCD ratios $R_{\eta_{c}/J/\psi}^{\pi}=1.74^{+0.66}_{-0.50}$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}S^{+})_{\Delta S=0}/{\rm BR}(B_{c}^{+}\to\eta_{c}S^{+})_{\Delta S=1}\sim{\cal O}(0.01)$ might be urgently demanded, which could help decipher the QCD dynamics involved in the considered $B_{c}$-meson decays, as well as in the related charmonia $\eta_{c}$ and $J/\psi$ greatly.
@@ Acknowledgements.
X.L. thanks Vanya Belyaev and Pei-Lian Li for their helpful communications. This work is supported by the National Natural Science Foundation of China under Grant No. 11875033 and by the Qing Lan Project of Jiangsu Province (No. 9212218405).

### v3

@@ III.2 $B_{c}^{+}\to\eta_{c}(A,S,T)^{+}$
To understand this peculiar feature in the $B_{c}^{+}\to\eta_{c}S^{+}$ decays, the related decay amplitudes calculated in the iPQCD formalism are presented explicitly in Table 4. For effective and clear clarifications, we also quote the available $B_{c}^{+}\to J/\psi S^{+}$ decay amplitudes Liu:2023kxr and present them in Table 4 simultaneously. Evidently, the originally constructive interferences to the $B_{c}^{+}\to J/\psi S^{+}$ BRs between the nonfactorizable-emission diagrams in Figs. 1(c) and 1(d) with antisymmetric leading-twist distribution amplitude of light scalars do not appear in the $B_{c}^{+}\to\eta_{c}S^{+}$ modes. In contrast, the sharply destructive interferences result in much smaller decay amplitudes as exhibited in the third column of Table 4. Furthermore, the destructions in $S1$ are heavier than those in $S2$ for the $B_{c}^{+}\to\eta_{c}(a_{0}(1450),K_{0}^{*}(1430))^{+}$ decays. The situation is very contrary to that in the $B_{c}^{+}\to J/\psi S^{+}$ decays Liu:2023kxr . Due to allowed $SU(3)$ symmetry-breaking effects, the comparably large factorizable-emission contributions, although evidently suppressed relative to those of $B_{c}^{+}\to\eta_{c}(P,V)^{+}$, lead to further significant destructions between the decay amplitudes of factorizable-emission and nonfactorizable-emission topologies in these $B_{c}^{+}\to\eta_{c}S^{+}$ decays. However, it should be stressed that, due to the light quark masses objectively satisfying the relation, that is, $m_{s}\gg m_{d}\sim m_{u}$, the apparently larger factorizable-emission contributions induced by the vector decay constants $f_{\kappa^{+}},f_{K_{0}^{*}(1430)^{+}}\propto(m_{s}-m_{u})$ Cheng:2005nb are therefore produced in $B_{c}^{+}\to\eta_{c}(\kappa,K_{0}^{*}(1430))^{+}$ channels, relative to the factorizable-emission amplitudes in $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$ modes proportional to the value of $(m_{d}-m_{u})$, namely, tiny broken isospin symmetry. The above iPQCD results predicted for the $B_{c}^{+}\to\eta_{c}S^{+}$ decays await the future experimental tests, which could help us to further explore the complicated QCD dynamics potentially.
@@
Lastly, we come to the discussions of $B_{c}$ decays to $\eta_{c}$ plus a light tensor state. For the $B_{c}^{+}\to\eta_{c}T^{+}$ modes, to our best understanding, tensor states cannot be produced via the vector current and the conventionally factorizable-emission contributions at leading order are thus vanished naturally. Therefore, their studies must go beyond naive factorization. Then, the $B_{c}^{+}\to\eta_{c}T^{+}$ BRs calculated in the iPQCD formalism are
@@
associated with the ratios
$\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})}{{\rm BR}(B_{c}^{+}\to J/\psi\pi^{+})}$ $=$ $0.11^{+0.05}_{-0.04}\;,\qquad R^{\eta_{c}}_{a_{2}/\pi}\equiv\frac{{\rm BR}(B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+})}{{\rm BR}(B_{c}^{+}\to\eta_{c}\pi^{+})}=0.07^{+0.00}_{-0.01}\;,$ (83)
where the dominant errors come from the $B_{c}$-meson distribution amplitude shape parameter $\beta_{B_{c}}$. Future tests of these values, which are solely from nonfactorizable-emission contributions, would help us examine the reliability of this iPQCD formalism.
Similar to $R_{K^{*}/\rho}^{\eta_{c}}$ in the $B_{c}^{+}\to\eta_{c}V^{+}$ sector, another ratio $R^{\eta_{c}}_{K_{2}/a_{2}}$ in the $B_{c}^{+}\to\eta_{c}T^{+}$ decays is also defined by utilizing the $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}$ and $B_{c}^{+}\to\eta_{c}{K_{2}^{*}}(1430)^{+}$ BRs in the iPQCD framework, and its value is then read as
@@
This result is induced by only nonfactorizable-emission contributions and is very close to the value presented in Eq. (50), as well as that naively anticipated in factorization ansatz for factorizable-emission predominated $B_{c}^{+}\to\eta_{c}(\rho,K^{*})^{+}$ modes within errors.
@@
Analogous to $B_{c}^{+}\to\eta_{c}b_{1}(1235)^{+}$ but different from $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$, the considerably constructive interferences due to the antisymmetric leading-twist distribution amplitude between the two nonfactorizable-emission diagrams in Figs. 1(c) and 1(d) work in the $B_{c}^{+}\to\eta_{c}a_{2}(1320)^{+}$ mode indeed. The future measurements on the iPQCD predictions of $B_{c}^{+}\to\eta_{c}L^{+}$ in this work could help test, even differentiate, the reliability of the adopted approaches in the literature, which might help us to further understand the rich QCD dynamics in $B_{c}$-meson decays through pulling together disparate ideas.
IV Conclusions
@@
In summary, we have studied the $B_{c}$-meson decays into $\eta_{c}$ plus a light charged meson in the self-consistent iPQCD formalism at leading order of strong coupling $\alpha_{s}$. The numerical results and phenomenological insights on CP-averaged BRs in association with relative ratios are presented explicitly. The $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ and $B_{c}^{+}\to\eta_{c}a_{1}(1260)^{+}$ BRs are generally larger than those $B_{c}\to J/\psi$ decay ones correspondingly in the iPQCD formalism, which reveal the remarkably different QCD dynamics between these two kinds of $B_{c}$-meson decays. For (near-)future tests at experiments, the BRs of multibody modes of $B_{c}^{+}\to\eta_{c}(\pi,\rho)^{+}$ via $\eta_{c}$ resonance through $\eta_{c}\to p\bar{p}$ and $\eta_{c}\to\pi^{+}\pi^{-}(\pi^{+}\pi^{-},K^{+}K^{-},p\bar{p})$ are also predicted under the narrow-width approximation. The experimental search performed with the successfully upgraded LHCb detector for the BRs around ${\cal O}(10^{-6})$ and above, for example, ${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p})\pi^{+})=(2.70^{+0.74}_{-0.59})\times 10^{-6}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}(\to p\bar{p})\rho^{+})=(7.14^{+2.03}_{-1.60})\times 10^{-6}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}[\to 2(\pi^{+}\pi^{-})]\pi^{+})=(1.95^{+0.59}_{-0.50})\times 10^{-5}$, ${\rm BR}(B_{c}^{+}\to\eta_{c}[\to 2(\pi^{+}\pi^{-})]\rho^{+})=(5.16^{+1.62}_{-1.34})\times 10^{-5}$, $\cdots$, are expected to help probe the relevant decay channels and explore the nature of charmonium $\eta_{c}$. The almost equal BRs of $B_{c}^{+}\to(J/\psi,\eta_{c})b_{1}(1235)^{+}$, while the surprisingly smaller ones of $B_{c}^{+}\to\eta_{c}a_{0}(980,1450)^{+}$ compared to those of $B_{c}^{+}\to J/\psi a_{0}(980,1450)^{+}$ need further explorations in both aspects of theory and experiment. Moreover, the predicted ratios $R_{K/\pi}^{\eta_{c}}$ and $R_{K^{*}/\rho}^{\eta_{c}}$ matching the anticipations based on the naive ansatz means that the $B_{c}^{+}\to\eta_{c}(P,V)^{+}$ decays are predominated by factorizable-emission contributions. The experimental tests on the iPQCD ratios $R_{\eta_{c}/J/\psi}^{\pi}=1.74^{+0.66}_{-0.50}$ and ${\rm BR}(B_{c}^{+}\to\eta_{c}S^{+})_{\Delta S=0}/{\rm BR}(B_{c}^{+}\to\eta_{c}S^{+})_{\Delta S=1}\sim{\cal O}(0.01)$ might be urgently demanded, which could help decipher the QCD dynamics involved in the considered $B_{c}$-meson decays as well as in the related charmonia $\eta_{c}$ and $J/\psi$ greatly.
@@ Acknowledgements.
X.L. thanks Vanya Belyaev and Pei-Lian Li for their helpful communications. This work is supported by the National Natural Science Foundation of China under Grant No. 11875033 and by the Qing Lan Project of Jiangsu Province (Grant No. 9212218405).